OpenStudy (dls):

Solve for x and y if xy=0 and x+y=1

jhonyy9 (jhonyy9):

do you have any idea ?

OpenStudy (dls):

umm I though of substituting x+y=1 x=(1-y) y(1-y)=0 but this gives y=0 as well as y=1.. x and y are 0 and 1 but which one is which idk.

Parth (parthkohli):

It seems obvious that the solutions are (0,1) and (1,0)

Parth (parthkohli):

Since the product is zero, then one of the numbers is zero.

OpenStudy (dls):

yes but this is a question from CO-ORDINATE and i have both the choices.. 0,1 and 1,0..

Parth (parthkohli):

Symmetrical equations.

OpenStudy (dls):

hein?

Parth (parthkohli):

If you have both the options, then select both.

OpenStudy (dls):

single choice

Parth (parthkohli):

Screw single choice.

OpenStudy (dls):

If I say x(x+y)=1 then

OpenStudy (dls):

o.O ?

OpenStudy (dls):

x=1 and y=0

Parth (parthkohli):

?